Given that P(X=2)=8P(X=4)+80P(X=46) for a Poisson distribution variate X. Find

1) P(X<2)

2) P(X>4)

3) P(X>=1)


1
Expert's answer
2022-03-03T10:01:46-0500

For Poisson's distribution, "P(X=x)"  is given by 


"P(X=x)=\dfrac{e^{-\lambda}\cdot\lambda^x}{x!}"


It is given that "P(X=2)=8P(X=4)+80P(X=6)"


"\dfrac{e^{-\lambda}\cdot\lambda^2}{2!}=8\cdot\dfrac{e^{-\lambda}\cdot\lambda^4}{4!}+80\cdot\dfrac{e^{-\lambda}\cdot\lambda^6}{6!}"

"\dfrac{\lambda^2}{2}=\dfrac{\lambda^4}{3}+\dfrac{\lambda^6}{9}"

Since "\lambda>0" we obtain


"2\lambda^4+6\lambda^2-9=0"

"\lambda^2=\dfrac{-3+3\sqrt{3}}{2}, \lambda>0"

"\lambda=\sqrt{\dfrac{-3+3\sqrt{3}}{2}}\approx1.0479"


1)


"P(X<2)=P(X=0)+P(X=1)"

"=\dfrac{e^{-1.0479}\cdot(1.0479)^0}{0!}+\dfrac{e^{-1.0479}\cdot(1.0479)^1}{1!}"

"\approx0.71814"

2)


"P(X>4)=1-P(X=0)-P(X=1)"

"-P(X=2)-P(X=3)="

"=1-\dfrac{e^{-1.0479}\cdot(1.0479)^0}{0!}-\dfrac{e^{-1.0479}\cdot(1.0479)^1}{1!}"

"-\dfrac{e^{-1.0479}\cdot(1.0479)^2}{2!}-\dfrac{e^{-1.0479}\cdot(1.0479)^3}{3!}"

"\approx0.00445"

3)


"P(X\ge1)=1-P(X=0)"

"=1-\dfrac{e^{-1.0479}\cdot(1.0479)^0}{0!}\approx0.64933"